OpenStudy (anonymous):

If v(t) = 2t + 1, the velocity of an object over time, what is the acceleration of the object at t = 1?

OpenStudy (anonymous):

@rajathsbhat can u help?

OpenStudy (anonymous):

do you know how to differentiate functions?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i don't see how else you could answer this question... i'll teach you some basics you need to solve this problem: 1: \(\Large \frac{d(t^{n})}{dt}=n*t^{n-1}\) 2: \(\Large \frac{d(k)}{dt}=0\) 3: \(\Large \frac{d(k * t^{n})}{dt}=k*n*t^{n-1}\) You should use these rules to differentiate the velocity because a=dv/dt. Tell me what you get.

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

so t = 1 what is d, k? and n?

OpenStudy (anonymous):

k is a constant, like 1 or 2 or whatever.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

n is the power to which t is raised. In this case, it's would be 1. and you should substitute t=1 only AFTER you've differentiated the function.

OpenStudy (anonymous):

*it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is d?

OpenStudy (anonymous):

it's called a differential. it's just a notation. it's looks like this: \[\large \frac{d}{dt} (whatever)\]

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

so will it be d(1^1)/d(1) = 1 x 1^-1

OpenStudy (anonymous):

no you only substitute for t after you're done with the diffentiation. so it's like \[\large \begin{align}\frac{d}{dt}(2t+1)&=\frac{d}{dt}(2t)+\frac{d}{dt}(1)\\\\ &=2\times1\times t^{1-1}+0\\\\ &=2.\end{align}\]

OpenStudy (anonymous):

not that you don't have t anymore. that means that the acceleration doesn't depend on the time. I.e. you have constant acceleration.

OpenStudy (anonymous):

IF you ended up with something like 2(t) in the end, you'd then substitute t=1 or 2 or anything that your question asks to get the acceleration at that time.